OpenStudy (anonymous):

Convert the decimal to a fraction. Be sure to show all of your work. 0.55

OpenStudy (anonymous):

55/100 divide both on 5 11/20

OpenStudy (asnaseer):

you can look at this in the following way:\[0.55=\frac{0.55}{1}=\frac{0.55}{1}\times\frac{100}{100}=\frac{0.55\times100}{100}=\frac{55}{100}=\frac{11}{20}\]

OpenStudy (anonymous):

thats what i got